OpenStudy (anonymous):

Simplify the expression below. -5-8

OpenStudy (ghazi):

-x-x=-2x similarly -5-8=-13

OpenStudy (ghazi):

@britbrat4601 clear?

OpenStudy (anonymous):

im really bad at math and i have thia pt test and i cant remeber anytghing!

OpenStudy (ghazi):

you'll go through it easily ....chill out

OpenStudy (anonymous):

well ya that one was common sense i guess lol thank u

OpenStudy (ghazi):

:) YW

jimthompson5910 (jim_thompson5910):

If you have two negatives, or you're subtracting a negative from a negative, then add the two numbers (ignore the negatives). Then make the final answer positive

jimthompson5910 (jim_thompson5910):

-5 - 8 Add: 5+8 = 13 Then make final answer negative: -13 So - 5 - 8 = -13